The energy required to remove an electron from M shell of H-atom is 1.51 eV, then the energy of the first excited state will be:
 
  1. -1.51 eV
  2. +1.51 eV
  3. -3.4 eV
  4. -13.6 eV

Grade:12th pass

1 Answers

Jotin Thingnam
6 years ago
Energy =-13.6/n2Where n= energy levelFirst excitation means, n=2Therefore, E= -13.6/4 eV = -3.4 eV